The total number of subtractions that result in 11111 remaining after a four-digit number has been subtracted form a five-digit numer and the digits 1 through 9 have all been used is ______?

PS. I was told that the answer is 24, but I don't really understand why... Thank you so much for helping!

Answer by Edwin McCravy(20056)


Yes it's 24.  Here is why:

ABCDE
-FGHI
-----
11111

It's clear that there can't be any "borrowing"
in the subtraction.

It's clear that A can only be 1.  

So B-F = C-G = D-H = E-I = 1 and:

{(B,F),(C,G),(D,H),(E,I)} c {(3,2),(4,3),(5,4),(6,5),(7,6),(8,7),(9,8)}

We can't use (8,7) because then we'd have to use (6,5), and (4,3), but
that wouldn't allow us to use 9 or 2.

So we can only use (9,8), (7,6), (5,4) and (3,2)  

Here is the smallest possibility:

13579
-2468
-----
11111

There are 4!=24 ways to rearrange the last 4 
columns that go after the first one. The
first column can only be:

1
_
1

Then there are 4!=24 ways we can arrange these 4
column after it:

3, 5, 7, 9
2, 4, 6, 8
-, -, -, -
1, 1, 1, 1
 
Edwin